OpenStudy (anonymous):

Show steps on how to find derivative of y = (x^2 - 2x)^(1/2)

OpenStudy (shamil98):

\[y = \sqrt{x^2 -2x}\] \[y' =\]

OpenStudy (ness9630):

Use the product rule, I think that will work.

OpenStudy (ness9630):

Take the derivative of the outside function, then multiply it by the inside

OpenStudy (ness9630):

So: \[\LARGE (x^2-2x)^{1/2}=\frac{1}{2(x^2-2x)}*2x-2\] Make some sense?

OpenStudy (ness9630):

Sorry that should be: \[\sqrt{x^2-2x}\]

OpenStudy (shamil98):

I understand where you get 2x-2 from, but why is there a fraction?

OpenStudy (ness9630):

Sham, catch up bro: \[\LARGE x^{1/2}=\frac{1}{2}*x^{-1/2}\]

OpenStudy (ness9630):

@shamil98 understand?

OpenStudy (ness9630):

@NicholasWong752 Do YOU understand?

OpenStudy (ranga):

Power rule followed by chain rule. Or, Let t = x^2 - 2x y = t^1/2 dy/dx = (dy/dt)(dt/dx) dy/dt = (1/2t^(-1/2) = 1/(2(t^1/2)) = 1/(2(x^2 - 2x)^1/2) dt/dx = 2x - 2 dy/dx = (2x - 2)/(2(x^2 - 2x)^1/2) = (x - 1)/(x^2 - 2x)^1/2

OpenStudy (ness9630):

Where are you guys getting that (x-1)?

OpenStudy (shamil98):

\[\sqrt{x^2-2x} = \frac{ 1 }{ 2(x^2 -2x) } \times 2x -2\] \[\sqrt{x^2-2x} = \frac{ (x-1) }{ (x^2-2x) } \]

OpenStudy (shamil98):

2(x-1)/2(x^2-2x) the twos cancel out :P

OpenStudy (anonymous):

I understand why there's fraction, but I got stuck at that point

OpenStudy (ness9630):

Oh, right I didn't simplify..

OpenStudy (shamil98):

x-1/√(x^2-2x) right?

OpenStudy (ness9630):

Need a total summary of everything we did? @NicholasWong752

OpenStudy (ness9630):

Yes sham, nice job kiddo